Fermi National Laboratory


Accelerator Update for February 21, 2005 - February 23, 2005

Monday February 21
The day shift began with Operations monitoring store 3997, and with Linac off due to a Klystron problem.

Operations reported at 8:19 AM that the Linac Klystron RF stations were back up and that the Antiproton Source (Pbar) resumed stacking. MiniBooNE also resumed taking beam.

The Neutron Therapy Facility (NTF) requested beam at 8:53 AM for patient treatment.

At 9:07 AM, Operations reported having problems with the Linac beam. The Linac problems started when NTF began taking beam. Operators contacted experts who tuned the line at 9:25 AM.

Operations put the Recycler into shot setup to stash antiprotons at 9:41 AM.

At 10:30 AM, Operations received reports of a LCW leak on a Lambertson power supply (LAM41) in Pbar. Experts investigated and reported that the leak was small and not near anything critical. A repair would take about four hours, so this was put off for later.

Operations reported that by 11:31 AM the Recycler had stashed seven transfers of antiprotons. Pbar resumed stacking.

At 3:54 PM, Operations reported a decrease in the Tevatron (TeV) colliding beam luminosity. Operators believed it was due to a TeV damper problem.

NuMI experts began a study period at 5:42 PM.

Operations terminated store 3997 at 7:13 PM.

Operations loaded 36x0 study-store to check the dampers at 8:10 PM.

Pbar resumed stacking at 8:19 PM.

Operations terminated the study-store at 9:09 PM and put the TeV into shot setup at 10:03 PM.

NuMI studies ended at 10:22 PM.

Tuesday February 22
The midnight shift began with the TeV shot setup on hold due to a power supply problem (V901), and with MiniBooNE taking beam.

Operations put the TeV through a dry squeeze at 2:20 AM.

A power supply expert reported at 3:58 AM that he had replaced an internal 24-volt power supply on V901.

Operation established store 4000 at 5:38 AM with and initial luminosity of 77.3E30. Pbar resumed stacking.

A NuMI expert began a study at 5:43 AM.

Operations resumed sending beam to the Meson experiment MCenter at 5:57 AM.

The Recycler began stashing antiprotons from Pbar around 11:30 AM. Pbar resumed stacking at noon.

TeV expert flew wires for a study at 1:02 PM.

Pbar experts began a study period at 1:18 PM.

MCenter resumed taking beam at 1:21 PM.

NuMI experts reported at 6:07 PM that their beam commissioning was done. They would next want beam on Friday, 2/25/05.

At 6:48 PM, Operations reported that MI beam permit had tripped a few times due to MECAR. The permit reset each time.

At 6:52 PM, Operations reported that a Linac quadrupole power supply (L:QPS130) was having problems with regulation. This power supply has had a number of problems looked at by experts all day.

At 7:02 PM, Operations reported that the TeV sector D3 wet engine alarmed on an over temperature indication. Cryo system technicians investigated and put a fan on the engine. They will repair it after the store ends.

Operations put the Recycler into shot setup to stash antiprotons at 7:47 PM. Shot setups disable Switchyard (SY) beam.

Operations completed the stash and put Pbar back into stacking mode at 9:15 PM.

Operations resumed sending beam to MCenter at 9:36 PM.

At 10 PM, Operations reported that MI had tripped due to MECAR again. Over the next hour and a half, it tripped off a number of times. An expert was notified.

At 10:37 PM, a Linac expert turned off L:QPS130 to replace its power supply. The work was completed by 11:21 PM. Beam to Pbar, MiniBooNE, and MCenter all resumed.

Wednesday February 23
The midnight shift began with Operations monitoring store 4000 and the Recycler stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 12:24 AM, MECAR caused MI to trip. It reset okay.

At 7:22 AM, two Linac quadrupole power supplies began not regulating properly. Operators and an expert investigated, but of course, it showed no problems.

Pbar resumed stacking at 7:43 AM.

The Plan for Wednesday
The plan for today is to allow TeV experts to conduct a study before the store termination that is scheduled for sometime around noon. The next store will go in later this afternoon.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 2/23/2005   email Fermilab